# Local Thermal Operations and Classical Communication

### v1

@@ 4 Detailed results on $\mathcal{S}$LTOCC
Similarly to the tri-stochastic tensors, the bi-thermal tensors are defined by a finite set of linear (in)equalities (20), and thus they form a convex polytope. Moreover, to fully describe this set it is sufficient to consider only a discrete set of extremal points. Let us start with mentioning the limit of infinite temperature, $\beta=0$ – tristochastic tensors. In this case, the problem of characterizing vertices was already extensively studied in [57], where they have been divided into the permutation and non-permutation extreme points.
@@
In the finite temperature case, $\beta>0$, the situation is more convoluted, even when restricted to standard thermal operations [29]. As presented in [58], the structure of vertices for the thermal equivalent of Birkhoff polytope depends on relations between systems energies and temperature. Thus we start the discussion by presenting a simple, yet comprehensive, example of a bi-thermal operation on a 2-level system.
@@
Note that for two level system the defining relations of bi-thermal operations (20), determine all tensor values as a function of only one. Thus geometrically it is equivalent to a line segment. To describe its extremal points let us denote the energies by $E_{0}<E_{1}$, moreover to simplify the notation we set $g_{1,0}=e^{-\beta(E_{1}-E_{0})}$. The extremal bi-thermal tensors are:
@@
$T_{ijk}^{(2)}=\sum_{l}S_{il}^{(\beta)}T_{ljk}^{(1)}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ T_{ijk}^{(2)}=\sum_{l}T_{ilk}^{(1)}S_{lj}^{(\beta)}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ T_{ijk}^{(2)}=\sum_{l}T_{ljl}^{(1)}S_{lk}^{(\beta)}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \text{where }\leavevmode\nobreak\ S^{(\beta)}=\left(\begin{matrix}1-g_{1,0}&1\\ g_{1,0}&0\end{matrix}\right)\leavevmode\nobreak\ .$ (23)
This interconnection between extremal tensors is standard for tri-stochastic tensors [57], however in finite temperature it results in one tensor being "more-thermalized" than the second one and multiple aprication of thermal swaps generally leads to non-extremal tensors.
Figure 4: Attainable states $\mathbf{r}=T(\mathbf{p},\mathbf{q})$ in dimension $d=2$ using LTOCC lie between the extreme surfaces generated by $T_{1}$ (red-orange) and $T_{2}$ (blue-cyan). Each distribution is characterized by its first coefficient, eg. $\mathbf{p}=(p,1-p)$, and the extremal values are denoted by shaded lines. Although fully symmetric for $\beta=0$, the symmetry is lost for finite temperatures, $\beta>0$.
4.1 Approach to extremal points for high temperatures
To tackle the general structure of bi-thermal operations’ set in dimensions $d>2$ we start by considering the scenario of large temperatures as a deviation from $\beta=0$. To study the form of extremal bi-thermal tensors one can consider its series expansion around tri-stochastic tensors:
@@ 4.1 Approach to extremal points for high temperatures
Thus in the first order, the conditions for bi-thermality (20) lead to:
@@
Moreover, we must ensure the bi-thermal tensor $T_{ijk}$ is non-negative. Thus treating $A_{ijk}^{(1)}$ as tangent vector in the point $T_{ijk}$, this requirement corresponds to:
@@
The condition (27) does not restrict the values of all tangent vector $A_{ijk}^{(1)}$ elements since not all directions from $\tilde{T}_{ijk}$ are prohibited. In particular $A_{ijk}^{(1)}$ can always point in the centre of set of bi-thermal tensor with arbitrary amplitude. However, we are interested in the extremal bi-thermal tensors. Thus we focus on extremal tensor $T_{ijk}$ and extremal tangent vectors $T_{ijk}^{(1)}$, which saturate as many of (27) inequalities as possible.
This approach not only lets us recover the first order of extremal tensors expansion in dimension $d=2$ (21),(22), but also find the first order expansion in dimension $d=3$ around an exemplar permutation tensor (11), which yields $67$. The solutions, together with the code used to obtain them are presented in the github repository [59].
We note the enormous growth of extremal tensors’ number in finite temperature, even in such a small local dimension. This indicates that the complexity of the set of bi-thermal operations is immense even in comparison with standard thermal operations, for which in local dimension $d=3$ there were no more than $3$ extremal operations corresponding to one permutation [58].
4.2 Zero temperature
The opposite extremal case of zero temperature $\beta\to\infty$ turns out to be surprisingly simple. As stated in [42], for temperatures low enough, if the thermal bath has any spectral gap, effectively its entire population lies in the ground state. In such a scenario, thermal operations reduce to cooling maps defined by condition $\Lambda_{ij}=0$ if $i>j$. Moreover, in the appendix B, we present an alternative derivation of cooling maps, as a zero temperature limit $\beta\to\infty$ of thermal operations.
Knowing the limit form of thermal operations in zero temperature, we can combine them into bi-thermal operations.
Theorem 3.
In zero temperature, assuming a non-degenerated Hamiltoninan spectrum, the set of bi-thermal operations simplifies to a polytope of stochastic matrices satisfying $T_{ijk}=0$ if $i>j$ or $i>k$. The extremal operations correspond to tensors with 00 and $1$ entries, restricted only by the abovestated condition.
Hereafter we will call the zero temperature version of bi-thermal operations the bi-cooling operations.
@@ 4.2 Zero temperature
From the discussion of thermal operations in the zero temperature $\beta\to\infty$, we deduce, that the bi-thermality conditions (20) changes to
$T_{ijk}=0\text{ if }i>j\text{ or }i>k\leavevmode\nobreak\ .$ (28)
Thus effectively the columns of the tensor $T_{ijk}$ are decoupled and can be treated as independent probability vectors subject only to the condition (28). From this, the form of extremal tensors, as a combination of extremal columns, follows immediately. ∎
We note, that the set defined by the limiting form of bi-thermality conditions (28), might be larger than the limit of the set of bi-thermal tensors. However, we choose the first one as bi-thermal tensors in zero temperature, since in the discussed framework the fundamental physical actions are thermal operations. In the limit of zero temperature, one can combine them into any tensors satisfying conditions (28).
4.3 Explicit construction of a family of extreme bi-thermal tensors in low temperatures.
Although the general construction of extremal bi-thermal tensor seems unattainable, based on the results introduced in [58] we may construct a family of extreme bithermal tensors, which exist for every Hamiltonian when the temperature is low enough. Hereafter, for the sake of clarity, we adopt the notation $g_{i,j}=e^{-\beta(E_{i}-E_{j})}$ [58]. In dimension $2$ our example is just an extremal tensor (21)
@@
$=\left(\begin{matrix}1-\sum_{i=1}^{n}g_{i,0}&\vline&\begin{matrix}1&\cdots&1\\ \end{matrix}\\ \hline\cr\begin{matrix}g_{1,0}\\ \vdots\\ g_{d,0}\\ \end{matrix}&\vline&\mbox{\Large 0}\end{matrix}\right.\leavevmode\nobreak\ \left|\leavevmode\nobreak\ \leavevmode\nobreak\ \begin{matrix}1&\vline&\begin{matrix}0&\cdots&0\\ \end{matrix}\\ \hline\cr\begin{matrix}0\\ \vdots\\ 0\\ \end{matrix}&\vline&T_{i,0,k}|_{d=n-1}\end{matrix}\leavevmode\nobreak\ \leavevmode\nobreak\ \right|\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \cdots\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \left|\leavevmode\nobreak\ \leavevmode\nobreak\ \begin{matrix}1&\vline&\begin{matrix}0&\cdots&0\\ \end{matrix}\\ \hline\cr\begin{matrix}0\\ \vdots\\ 0\\ \end{matrix}&\vline&T_{i,n-1,k}|_{d=n-1}\end{matrix}\right)\leavevmode\nobreak\ .$
@@
$\sum_{i=1}^{n}g_{i,0}\leq 1\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{i=2}^{n}g_{i,1}\leq 1\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \cdots,\leavevmode\nobreak\ g_{n,n-2}+g_{n-1,n}\leq 1\leavevmode\nobreak\ .$ (32)
Each of those equations introduces some upper bound for the temperature, thus it is sufficient to satisfy the tightest of them. The above bi-thermal tensor is symmetric in the sense $T_{ijk}=T_{ikj}$, see fig. 5, thus the opposite way of slicing it would give the same formulas.
@@
Furthermore, each layer cannot be decomposed as a product of two other thermal operations, as demonstrated in [58], so this tensor cannot be obtained from anything "simpler" using local thermal pre- and post-processing. Finally in the limit of zero temperature, the bi-thermal tensor (31) simplifies to a bi-cooling tensor of least possible cooling i.e.
@@
Figure 5: Graphical representation of extremal bi-thermal tensor (31) in dimension $d=4$ as a cube and its separation into consecutive slices. Grey cells correspond to value $1$, red-to-blue to other other nonzero values, and empty cells to 00.
4.4 Local action
Having discussed the general structure of the set of bithermal operations, including methods for obtaining the extreme points, we shift our attention to the images of pairs of states under $\mathcal{S}$LTOCC. We start our discussion by considering local results. In other words, we will consider what Alice can obtain in her subsystem, excluding the created correlations with Bob’s subsystem. Thus, we will focus on one (bi-)thermal tensor, rather than the entire scheme. We stress that although the entire sub-section is presented from the perspective of (bi-)thermal tensors, the presented result holds also for tristochastic tensors.
It turns out that there is a strong connection between the images of (bi-)thermal tensors and thermomajorization. In the following theorem in square brackets, we present its extended version for bi-thermal tensors.
Theorem 4.
Let $T$ be a [bi-]thermal tensor, $\forall_{q}$ $T({\mathbf{\gamma}},\mathbf{q})={\mathbf{\gamma}}$ [and $T({\mathbf{q},\mathbf{\gamma}})={\mathbf{\gamma}}$], if and only if for any input distrstibutions $p,q$ the output is a probability distribution satisfying thermomajorization relations $\mathbf{p}\succ_{(\beta)}T(\mathbf{p},\mathbf{q})$ [and $\mathbf{q}\succ_{(\beta)}T(\mathbf{p},\mathbf{q})$].
@@ 4.4 Local action
In the case of bi-thermal tensors, one only needs to repeat the above arguments replacing the first with the second input state. ∎
As one can see, the problem of majorization between inputs and outputs is very similar to the scenario with thermal operations. On the other hand, the problem of finding a tensor, that may transform a pair of distributions into a desired one is much more complicated. The majorization of possible output by inputs is only the necessary conditions for possible mapping.
Lemma 1.
Let $\mathbf{p},\mathbf{q},\mathbf{r}$ be three probability distributions, then there exists a bi-thermal tensor $T$ mapping $\mathbf{p},\mathbf{q}$ into $\mathbf{r}$: $\mathbf{r}=T(\mathbf{p},\mathbf{q})$ if and only if $\mathbf{r}$ is majorized by a convex combination of majorization curves of $T^{(i)}(\mathbf{p},\mathbf{q})$ for all extremal tensors $T^{(i)}$.
@@
Suppose some combination of majorization curves for $T^{(i)}(\mathbf{p},\mathbf{q})$ majorizes $\mathbf{r}$. In that case, one can perform a convex combination with the same weights between appropriate extremal tensors $T^{(i)}$, obtaining tensor $\tilde{T}$ such that $\tilde{T}(\mathbf{p},\mathbf{q})\succ_{(\beta)}\mathbf{r}$. Thus additional thermal post-processing is sufficient to construct the desired tensor. Note that such post-processing can be included in bi-thermal tensor without changing its defining properties.
@@
From the above lemma, we can deduce the following observations, which simplify the discrimination of possible transforming bi-thermal tensors.
Observation 1.
Let $l$ be a curve obtained by, at each point, taking a maximum over majorization curves for $T^{(i)}(\mathbf{p},\mathbf{q})$, with $T^{(i)}$ being once again an extremal tensors. If $l$ does not majorize the curve corresponding to $\mathbf{r}$, then there exists no tristochastic (bi-thermal stochastic) tensor $T$, such that $T(\mathbf{p},\mathbf{q})=\mathbf{r}$.
Observation 2.
@@
The heuristic argument in favour of Observation 2 is as follows. The only tensors that map all pairs of basis vectors $e_{i},e_{j}$ into basis vectors are permutation tensors. Thus for any pair of input basis vectors its image under any permutation tensor majorizes its image under any tristochastic tensor. Since all finite probability distributions are convex combinations of basis vectors, we expect that the dominance of permutation tensors holds in the general case.
Fortunately, there exists also a stronger necessary condition for the existence of bi-thermal tensor $T(\mathbf{p},\mathbf{q})\ =\mathbf{r}$ for given distributions $\mathbf{p},\mathbf{q},\mathbf{r}$. To present it let us introduce the following notion.
For any distribution $p$ let us define its projection $\tilde{\mathbf{p}}$ onto the boundary of the probability simplex, which we will refer to as boundary distribution of $\mathbf{p}$,
@@
Theorem 5.
@@
Then the necessary condition for $\mathbf{r}$ to be achivable by a bi-thermal tensor form $\mathbf{p},\mathbf{q}$ is not only $\mathbf{p}\succ_{\beta}\mathbf{r}$ and $\mathbf{q}\succ_{\beta}\mathbf{r}$ but also $\tilde{\mathbf{p}}\succ_{\beta}\overline{\mathbf{r}}$ and $\tilde{\mathbf{q}}\succ_{\beta}\overline{\mathbf{r}}$.
@@
Assume that there exists some bi-thermal tensor such that $T(\mathbf{p},\mathbf{q})=\mathbf{r}$. Then by the definition of boundary distributions:
@@
Thus we effectively boosted the requirements proceeding from thermomajorization relations, leveraging the fact, that the allowed operations are bi-thermal.

### v2

@@ 5 Basic building blocks – thermal and bithermal tensors
Similarly to the tri-stochastic tensors, the bithermal tensors are defined by a finite set of linear (in)equalities (25), and thus they form a convex polytope. Moreover, to fully describe this set it is sufficient to consider only a discrete set of extremal points. Let us start by mentioning the limit of infinite temperature, $\beta=0$ – tristochastic tensors. In this case, the problem of characterizing vertices was already extensively studied in [63], where they have been divided into the permutation and non-permutation extreme points.
@@
In the finite temperature case, $\beta>0$, the situation is more convoluted, even when restricted to standard thermal operations [29]. As presented in [62], the structure of vertices for the thermal equivalent of Birkhoff polytope depends on the relations between system’s energies and temperature. Thus we start the discussion by presenting a simple, yet comprehensive, example of a bithermal operation on a 2-level system.
@@
Note that for a two-level systems the defining relations of bithermal operations (25), determine all tensor values as a function of only one of them. Thus geometrically it is equivalent to a line segment. To describe its extremal points let us denote the energies by $E_{0}<E_{1}$, moreover to simplify the notation we set $g_{1,0}=e^{-\beta(E_{1}-E_{0})}$. The extremal bithermal tensors are:
@@
$T_{ijk}^{(2)}=\sum_{l}S_{il}^{(\beta)}T_{ljk}^{(1)}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ T_{ijk}^{(2)}=\sum_{l}T_{ilk}^{(1)}S_{lj}^{(\beta)}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ T_{ijk}^{(2)}=\sum_{l}T_{ljl}^{(1)}S_{lk}^{(\beta)}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak$ (29)
where $\text{where }\leavevmode\nobreak\ S^{(\beta)}=\left(\begin{matrix}1-g_{1,0}&1\\ g_{1,0}&0\end{matrix}\right)\leavevmode\nobreak\ .$
This interconnection between extremal tensors is standard for tri-stochastic tensors [63], however in finite temperature it results in one tensor being “more thermalized” than the second one and multiple applications of thermal swaps generally leads to non-extremal tensors.
Before discussing larger local dimensions we would like to contrast the presented structure of bithermal tensors with thermal tensors for local dimension $2$. Following [62] we note that in this dimension there are two extremal thermal operations identity $\mathbb{I}$ and thermal swap $S^{(\beta)}$ (29). Thus each of the two layers of extremal thermal tensor $T$ can be either extremal thermal operations giving four, not two, (linearly independent) extremal operations.
Figure 6: Attainable states $\mathbf{r}=T(\mathbf{p},\mathbf{q})$ in dimension $d=2$ using LTOCC lie between the extreme surfaces generated by $T_{1}$ (red-orange) and $T_{2}$ (blue-cyan). Each distribution is characterized by its first coefficient, e.g. $\mathbf{p}=(p,1-p)$, and the extremal values are denoted by shaded lines. Although fully symmetric for $\beta=0$, the symmetry is lost for finite temperatures, $\beta>0$.
5.1 Extremal points of thermal “Birkhoff” polytope in infinite and large temperatures
We start this section by recalling thermal and bithermal tensors in the limit of infinite temperatures $\beta\to 0$. In this case the condition of thermality (18) simplifies to:
$\forall_{k}\sum_{j}T_{ijk}\eta_{j}=\eta_{i}\iff\forall_{k}\sum_{j}T_{ijk}=1$
where $\eta_{i}=1/d$ is a flat state. Thus the layers of thermal tensors reduce to bi-stochastic matrices. Therefore, extremal stochastic tensors are the ones with permutation matrices as their layers since those are the extremal bi-stochastic matrices.
Similarly, for bithermal tensors, their defining conditions (25) in the infinite temperature limit reduces to the conditions for tristochastic tensors (10). Thus the extremal tensors in infinite temperature are, among others, permutation tensors.
To tackle the general structure of bithermal operations’ set in dimensions $d>2$ we start by considering the scenario of large temperatures as a perturbation from the infinite temperature $\beta=0$. To study the form of extremal bithermal tensors one can consider its series expansion around tri-stochastic tensors:
@@ 5.1 Extremal points of thermal “Birkhoff” polytope in infinite and large temperatures
Thus in the first order, the conditions for bithermality (25) lead to:
@@
Moreover, we must ensure the bithermal tensor $T_{ijk}$ is non-negative. Thus treating $A_{ijk}^{(1)}$ as tangent vector in the point $T_{ijk}$, this requirement corresponds to:
@@
The condition (33) does not restrict the values of all tangent vector $A_{ijk}^{(1)}$ elements since not all directions from $\tilde{T}_{ijk}$ are prohibited. In particular $A_{ijk}^{(1)}$ can always point in the centre of the set of bithermal tensors with arbitrary amplitude. However, we are interested in the extremal bithermal tensors. Thus we focus on extremal tensor $T_{ijk}$ and extremal tangent vectors $T_{ijk}^{(1)}$, which saturate as many of inequalities (33) as possible.
This approach not only lets us recover the first order of extremal tensors expansion in local dimension $d=2$ (27),(28), but also find the first order expansion in dimension $d=3$ around an exemplar permutation tensor (11), which yields $67$ distinct approximants. The solutions, together with the code used to obtain them are presented in the GitHub repository [64].
We note the enormous growth of extremal tensors’ number in finite temperature, even in such a small local dimension. This indicates that the complexity of the set of bithermal operations is immense even in comparison with standard thermal operations, for which in local dimension $d=3$ there were no more than $3$ extremal operations corresponding to one permutation [62].
5.2 Bithermal tensors approaching zero temperature
The opposite extreme of zero temperature $\beta\to\infty$ turns out to be surprisingly simple. As stated in [45], for temperatures low enough, if the thermal bath has any spectral gap, effectively its entire population lies in the ground state. In such a scenario, thermal operations reduce to cooling maps defined by condition $\Lambda_{ij}=0$ if $i>j$. Moreover, in the B, we present an alternative derivation of cooling maps, as a zero temperature limit $\beta\to\infty$ of thermal operations.
Knowing the limit form of thermal operations at zero temperature, we can combine them into thermal and bithermal operations.
Observation 3 (Extremal bicooling tensors).
In zero temperature, assuming a non-degenerated Hamiltonian spectrum, the set of thermal and bithermal tensors simplify to polytopes of stochastic matrices satisfying respectively $T_{ijk}=0$ if $i>j$ and
$T_{ijk}=0\text{ if }i>j\text{ or }i>k$ (34)
The extremal thermal and bithermal tensors are the ones with only 00 and $1$ entries, restricted only by the abovestated condition.
Hereafter we will call the zero temperature version of bithermal operations the bicooling operations.
@@ 5.2 Bithermal tensors approaching zero temperature
For the thermal requirement in the zero temperature $\beta\to\infty$, we have $T_{ijk}=0\text{ if }i>j$ and $T_{ijk}=0\text{ if }i>k$, thus (34) follows trivialy. From the condition (34) we see that the columns of the (bi-)cooling tensor are decoupled and can be treated as independent probability vectors subject only to the condition (34). Thus the form of extremal tensors, as a combination of extremal columns, follows immediately. ∎
We note, that the set defined by the limiting form of bithermality conditions (34) might be larger than the zero-temperature limit of the set of bithermal tensors. However, we choose the first one as bicooling tensors, since in the discussed framework the fundamental physical operations are thermal operations. In the limit of zero temperature, one can combine them into any tensors satisfying conditions (34).
5.3 Explicit construction of a family of extreme bithermal tensors in low temperatures
Although the general construction of extremal bithermal tensors seems unattainable, based on the results introduced in [62] we may construct a family of extreme bithermal tensors, which exist for every Hamiltonian when the temperature is low enough. Hereafter, for the sake of clarity, we adopt the notation $g_{i,j}=e^{-\beta(E_{i}-E_{j})}$ [62]. In dimension $2$ our example is just an extremal tensor (27)
@@ 5.3 Explicit construction of a family of extreme bithermal tensors in low temperatures
$=\left(\begin{matrix}1-\sum_{i=1}^{n}g_{i,0}&\vline&\begin{matrix}1&\cdots&1\\ \end{matrix}\\ \hline\cr\begin{matrix}g_{1,0}\\ \vdots\\ g_{d,0}\\ \end{matrix}&\vline&\mbox{\Large 0}\end{matrix}\right.\leavevmode\nobreak\ \left|\leavevmode\nobreak\ \leavevmode\nobreak\ \begin{matrix}1&\vline&\begin{matrix}0&\cdots&0\\ \end{matrix}\\ \hline\cr\begin{matrix}0\\ \vdots\\ 0\\ \end{matrix}&\vline&T_{i,0,k}|_{d=n-1}\end{matrix}\leavevmode\nobreak\ \leavevmode\nobreak\ \right|\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \cdots\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \left|\leavevmode\nobreak\ \leavevmode\nobreak\ \begin{matrix}1&\vline&\begin{matrix}0&\cdots&0\\ \end{matrix}\\ \hline\cr\begin{matrix}0\\ \vdots\\ 0\\ \end{matrix}&\vline&T_{i,n-1,k}|_{d=n-1}\end{matrix}\right)\leavevmode\nobreak\ .$
@@
$\sum_{i=1}^{n}g_{i,0}\leq 1\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{i=2}^{n}g_{i,1}\leq 1\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \cdots,\leavevmode\nobreak\ g_{n-1,n-2}+g_{n,n-2}\leq 1\leavevmode\nobreak\ .$ (38)
Each of those equations introduces some upper bound for the temperature, thus it is sufficient to satisfy the tightest of them. The above bithermal tensor is symmetric in the sense $T_{ijk}=T_{ikj}$, see fig. 7, thus the opposite way of slicing it would give the same formulas.
@@
Furthermore, each layer cannot be decomposed as a product of two other thermal operations, as demonstrated in [62], so this tensor cannot be obtained from anything “simpler” using local thermal pre- and post-processing. Finally, in the limit of zero temperature, the bithermal tensor (37) simplifies to a bi-cooling tensor of least possible cooling i.e.
@@
Figure 7: Graphical representation of extremal bithermal tensor (37) in dimension $d=4$ as a cube and its separation into consecutive slices. Grey cells correspond to value $1$, red-to-blue cells to other nonzero values, and empty cells to 00.
5.4 Majorization-like restrictions on outputs of (bi)thermal tensors
Having discussed the general structure of the set of thermal and bithermal tensors, including methods for obtaining the extreme points, we shift our attention to the images of pairs of states under LTOCC and $\mathcal{S}$LTOCC. We start our discussion by considering local results. In other words, we will consider what Bob can obtain in his subsystem after reciving information from Alice, excluding the correlations between subsystems. Thus, we will focus on one (bi)thermal tensor, rather than the entire scheme. We stress that although the entire sub-section is presented from the perspective of thermal operations, the presented results hold also for (tri)stochastic tensors.
It turns out that there is a strong connection between the images of (bi)thermal tensors and thermomajorization. In the following theorem in square brackets, we present its extended version for bithermal tensors.
Theorem 10 ((Bi)thermality from thermomajorisation).
The tensor $T$ is a [bi-]thermal tensor, $\forall_{q}$ $T({\mathbf{\gamma}},\mathbf{q})={\mathbf{\gamma}}$ [and $T({\mathbf{q},\mathbf{\gamma}})={\mathbf{\gamma}}$], if and only if for any input distributions $p,q$ the output is a probability distribution satisfying thermomajorization relations $\mathbf{p}\succ_{(\beta)}T(\mathbf{p},\mathbf{q})$ [and $\mathbf{q}\succ_{(\beta)}T(\mathbf{p},\mathbf{q})$].
@@ 5.4 Majorization-like restrictions on outputs of (bi)thermal tensors
In the case of bithermal tensors, one only needs to repeat the above arguments replacing the first with the second input state. ∎
As one can see, the problem of majorization between inputs and outputs is very similar to the scenario with thermal operations. For the thermal tensors, one can also provide an analogue of Theorem 2:
Theorem 11.
There exists a thermal tensor $T$, $T(\mathbf{\gamma},\mathbf{q})=\mathbf{\gamma}$, mapping $(\mathbf{p},\mathbf{q})$ to $\mathbf{r}$ if and only if $\mathbf{p}\succ_{\beta}\mathbf{r}$.
@@
The implication in the first direction follows directly from the above Theorem 10. The implication in the other follows from the fact that if $\mathbf{p}\succ_{\beta}\mathbf{r}$, then, by the Theorem 2, there exists a Gibbs-preserving stochastic matrix $\Lambda$ such that $\Lambda(\mathbf{p})=\mathbf{r}$. Thus we may construct a thermal tensor $T$ with all layers equal to $\Lambda$: $T_{ikl}=\Lambda_{ik}$ to obtain: $T(\mathbf{p},\mathbf{q})_{i}=\sum_{kl}\Lambda_{ik}p_{k}q_{l}=\sum_{k}\Lambda_{ik}p_{k}=r_{i}$, which ends the proof. ∎
On the other hand, the problem of finding a bithermal tensor, that may transform a pair of distributions into a desired one is much more complicated. The majorization of possible output by inputs is only a necessary condition for possible mapping.
Theorem 12.
Let $\mathbf{p},\mathbf{q},\mathbf{r}$ be three probability distributions, then there exists a bithermal tensor $T$ mapping $\mathbf{p},\mathbf{q}$ into $\mathbf{r}$: $\mathbf{r}=T(\mathbf{p},\mathbf{q})$ if and only if $\mathbf{r}$ is majorized by a convex combination of majorization curves of $T^{(i)}(\mathbf{p},\mathbf{q})$ for all extremal bithermal tensors $T^{(i)}$.
Proof.
Suppose some combination of majorization curves for $T^{(i)}(\mathbf{p},\mathbf{q})$ majorizes $\mathbf{r}$. In that case, one can perform a convex combination with the same weights between appropriate extremal tensors $T^{(i)}$, obtaining tensor $\tilde{T}$ such that $\tilde{T}(\mathbf{p},\mathbf{q})\succ_{(\beta)}\mathbf{r}$. Thus additional thermal post-processing is sufficient to construct the desired tensor. Note that such post-processing can be included in bithermal tensor without changing its defining properties.
@@
From the above lemma, we can deduce the following observations, which simplify the discrimination of possible transforming bithermal tensors.
Observation 4.
Let $l$ be a curve obtained by, at each point, taking a maximum over majorization curves for $T^{(i)}(\mathbf{p},\mathbf{q})$, with $T^{(i)}$ being once again extremal bithermal tensors. If $l$ does not majorize the curve corresponding to $\mathbf{r}$, then there exists no bithermal tensor $T$, such that $T(\mathbf{p},\mathbf{q})=\mathbf{r}$.
Observation 5.
@@
Note that Observation 5 is only conjectured, based on extensive numerical experiments. However, the heuristic argument in its favour can be given as follows. The only tensors that map all pairs of basis vectors $e_{i},e_{j}$ into basis vectors are permutation tensors. Thus for any pair of input basis vectors its image under any permutation tensor majorizes its image under any tristochastic tensor. Since all finite probability distributions are convex combinations of basis vectors, we expect that the dominance of permutation tensors holds in the general case.
Fortunately, there also exists a stronger necessary condition for the existence of bithermal tensor $T(\mathbf{p},\mathbf{q})\ =\mathbf{r}$ for given distributions $\mathbf{p},\mathbf{q},\mathbf{r}$. To present it let us introduce the following notion.
For any distribution $\mathbf{p}$ let us define its projection $\tilde{\mathbf{p}}$ onto the boundary of the probability simplex, which we will refer to as boundary distribution of $\mathbf{p}$,
@@
Theorem 13 (Enhanced thermomajorization for bithermal tensors).
@@
Then the necessary condition for $\mathbf{r}$ to be achivable by a bithermal tensor from $\mathbf{p},\mathbf{q}$ is not only $\mathbf{p}\succ_{\beta}\mathbf{r}$ and $\mathbf{q}\succ_{\beta}\mathbf{r}$ but also $\tilde{\mathbf{p}}\succ_{\beta}\overline{\mathbf{r}}$ and $\tilde{\mathbf{q}}\succ_{\beta}\overline{\mathbf{r}}$.
@@
Assume that there exists some bithermal tensor such that $T(\mathbf{p},\mathbf{q})=\mathbf{r}$. Then by the definition of boundary distributions:
@@
Thus we effectively boosted the requirements proceeding from thermomajorization relations, leveraging the fact, that the allowed operations are bithermal.